# Measurements of low energy nuclear recoil quenching factors for Na and I recoils in the NaI(Tl) scintillator

### v2

@@ V Results and discussions
QFs for sodium recoils, spanning a broad energy range from 3.8 to 151.6 keVnr, and iodine recoils ranging from 6.1 to 74.9 keVnr, have been successfully obtained, extending this investigation into lower energy regions. Leveraging the high-light-yield crystal, the achievement includes a remarkable measurement of 3.8 keVnr, representing the lowest nuclear recoil measurement for sodium to date. Through a thorough revisitation of previously reported data, involving updates in neutron energy calibration, event selections, as well as selection and trigger efficiency evaluations, the revisited QFs are presented alongside the new measurements, as depicted in Fig.9 (a). These results are compared with measurements from other research groups [26, 43, 25], showing consistent values. The QF values for each recoil angle obtained in this study are detailed in Tables 1 and 2.
The QFs presented in Fig.9(a) from various experiments have been linearly calibrated using either a 59.54 keV line or a 57.6 keV line. However, discrepancies in QF values arise from the utilization of different energy calibration methods involving distinct energy points, as extensively discussed in Ref. [30]. This discrepancy is primarily attributed to the non-proportionality of NaI(Tl) crystals [44, 45]. To address non-proportionality and ensure an accurate conversion between the measured charge (light output) and deposited energy, it becomes imperative to implement an energy-dependent calibration for both s and electrons. Our investigation into the non-proportionality phenomenon involved a thorough examination utilizing X-rays and -rays sourced from external radioactive materials, as well as internal or cosmogenic contaminants present in the COSINE-100 crystals [46]. This comprehensive study yielded an energy-dependent calibration curve, enabling the recalculation of QF results. Figure 9(b) shows the QF values obtained in this study, accompanied by the recalculated values considering the non-proportionality of the NaI(Tl) crystals for both and X-rays. These data points are modeled using the modified Lindhard model [47] as described in Ref. [11]. The inclusion of non-proportionality in the QF assessment reveals a notable increase in QF values, particularly for the low energy region, underscoring the importance of using appropriate QF values depending on the electron-equivalent energy calibration.
@@ VI Conclusion
Measurements of low energy QFs for NaI(Tl) crystals were conducted using a high light yield crystal of 26.0 0.7 NPE/keVee, achieved through a novel crystal encapsulation technique. The precise evaluation of QFs for low energy nuclear recoils was facilitated by an in-depth understanding of low energy scintillation events from waveform simulations. These measurements extended to sodium QFs as low as 3.8 keVnr, marking the first measurement for sodium below 5 keVnr. Furthermore, a revisitation of previously reported data, with an enhanced comprehension of incident neutron energy and a refined understanding of low energy scintillation events through waveform simulation, was reported. The measurements show general consistency with recent reports from other research groups. These low energy QF measurements for NaI(Tl) crystals hold relevance for future data analyses in the pursuit of dark matter and Coherent Elastic Neutrino-Nucleus Scattering (CENS) studies with NaI(Tl) crystals, such as the COSINE-100 and NEON experiments.
@@ Acknowledgements.
We thank the IBS Research Solution Center (RSC) for providing high-performance computing resources. This work is supported by the Institute for Basic Science (IBS) under project code IBS-R016-A1.References
@@
- [16] J. Xu, F. Calaprice, F. Froborg, E. Shields, and B. Suerfu, (SABRE Collaboration), SABRE – A test of DAMA with high-purity NaI(Tl) crystals, AIP Conf. Proc. 1672, 040001 (2015).
@@
- [30] D. Cintas et al., Quenching Factor consistency across several NaI(Tl) crystals, J. Phys. Conf. Ser. 2156, 012065 (2021).
@@
- [33] J. J. Choi et al., Waveform Simulation for Scintillation Characteristics of NaI(Tl) Crystal, arXiv:2402.17125.
@@
- [46] S. M. Lee et al., (COSINE-100 Collaboration), Nonproportionality of NaI(Tl) Scintillation Detector for Dark Matter Search Experiments, arXiv:2401.07462.

### v3

@@ V Results and discussions
QFs for sodium recoils, spanning a broad energy range from 3.8 to 151.6 keVnr, and iodine recoils ranging from 6.1 to 74.9 keVnr, have been successfully obtained, extending this investigation into lower energy regions. Leveraging the high-light-yield crystal, the achievement includes a remarkable measurement of 3.8 keVnr, representing the lowest nuclear recoil measurement for sodium to date. Through a re-examination of previously reported data, involving updates in neutron energy calibration, event selections, as well as selection and trigger efficiency evaluations, the reexamined QFs are presented alongside the new measurements, as depicted in Fig. 9(a). These results are compared with measurements from other research groups [26, 46, 25], showing consistent values. The QF values for each recoil angle obtained in this study are detailed in Tables 1 and 2.
The QFs presented in Fig.9(a) from various experiments have been linearly calibrated using either a 59.54 keV line or a 57.6 keV line. However, discrepancies in QF values arise from the utilization of different energy calibration methods involving distinct energy points, as extensively discussed in Ref. [30]. This discrepancy is primarily attributed to the nonproportionality of NaI(Tl) crystals [47, 48]. To address nonproportionality and ensure an accurate conversion between the measured charge (light output) and deposited energy, it becomes imperative to implement an energy-dependent calibration for both s and electrons. Our investigation into the nonproportionality phenomenon involved a thorough examination utilizing x rays and rays sourced from external radioactive materials, as well as internal or cosmogenic contaminants present in the COSINE-100 crystals [49]. This comprehensive study yielded an energy-dependent calibration curve, enabling the recalculation of QF results. Figure 9(b) shows the QF values obtained in this study, accompanied by the recalculated values considering the nonproportionality of the NaI(Tl) crystals for both and x rays. These data points are modeled using the modified Lindhard model [50] as described in Ref. [11]. The inclusion of nonproportionality in the QF assessment reveals a notable increase in QF values, particularly for the low energy region, underscoring the importance of using appropriate QF values depending on the electron-equivalent energy calibration.
@@ VI Conclusion
Measurements of low energy QFs for NaI(Tl) crystals were conducted using a high light yield crystal of 26.0 0.7 NPE/keVee, achieved through a novel crystal encapsulation technique. The precise evaluation of QFs for low energy nuclear recoils was facilitated by an in-depth understanding of low energy scintillation events from waveform simulations. These measurements extended to sodium QFs as low as 3.8 keVnr, marking the first measurement for sodium below 5 keVnr. Furthermore, a reexamination of previously reported data, with an enhanced comprehension of incident neutron energy and a refined understanding of low energy scintillation events through waveform simulation, was reported. The measurements show general consistency with recent reports from other research groups. These low energy QF measurements for NaI(Tl) crystals hold relevance for future data analyses in the pursuit of dark matter and coherent elastic neutrino-nucleus scattering (CENS) studies with NaI(Tl) crystals, such as the COSINE-100 and NEON experiments.
@@ Acknowledgements.
We thank the IBS Research Solution Center (RSC) for providing high-performance computing resources. This work is supported by the Institute for Basic Science (IBS) under project code no. IBS-R016-A1.References
@@
- [16] M. Antonello et al., Characterization of SABRE crystal NaI-33 with direct underground counting, Eur. Phys. J. C 81, 299 (2021).
@@
- [30] D. Cintas et al., A measurement of the sodium and iodine scintillation quenching factors across multiple NaI(Tl) detectors to identify systematics, arXiv:2402.12480.
@@
- [33] J. J. Choi et al., Waveform simulation for scintillation characteristics of NaI(Tl) crystal, Nucl. Instrum. Meth. A 1065, 169489 (2024).
@@
- [35] M. Manolopoulou, M. Fragopoulou, S. Stoulos, E. Vagena, W. Westmeier, and M. Zamani, Neutron spectrometry with He-3 proportional counters, Journal of Physics: Conference Series 366, 012033 (may, 2012).
- [36] D. Thomas, Neutron spectrometry, Radiation Measurements 45, 1178–1185 (2010).
@@
- [45] J. I. Collar, Light WIMP Searches: The Effect of the Uncertainty in Recoil Energy Scale and Quenching Factor, arXiv:1010.5187.
@@
- [49] S. M. Lee et al., (COSINE-100), Nonproportionality of NaI(Tl) scintillation detector for dark matter search experiments, Eur. Phys. J. C 84, 484 (2024).